OpenStudy (anonymous):

find the probability of selecting a class that runs greater than 51.5, uniformly distributed between 46.0 and 56.0

OpenStudy (anonymous):

Let \(X\) the random variable for this situation. Then the probability density is given by \[f_X(x)=\begin{cases}\dfrac{1}{56-46}=\dfrac{1}{10}&\text{for }46\le x\le56\\\\0&\text{elsewhere}\end{cases}\] Desired probability: \[P(X>51.5)=\int_{51.5}^{56}f(x)~dx=\frac{1}{10}(56-51.5)=0.45\]